OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line?

OpenStudy (anonymous):

4/5**

Nnesha (nnesha):

same like last one y = mx + b here m is slope and b is y-intercept

Nnesha (nnesha):

just replace m and by by their values

OpenStudy (anonymous):

I got y=4/5x+1

Nnesha (nnesha):

and that is right

OpenStudy (anonymous):

4y − 5x = 5 y − 4x = 5 5y + 4x = 5 5y − 4x = 5 those are my choices

Nnesha (nnesha):

they are in standard form

Nnesha (nnesha):

so do back word

Nnesha (nnesha):

can you solve first equation for y

OpenStudy (thecatman):

5y=5-4x and divide all numbers by 5

OpenStudy (anonymous):

I don't know how, @Nnesha

Nnesha (nnesha):

okay so for example first one  first add 5x both side